OpenStudy (anonymous):

The cubed root of 8x3y6z4 (the 3, 6, and 4 are powers)

OpenStudy (anonymous):

the cubed root of 8 is 2 the cubed root of \(x^3\) is \(x\) the cubed root of \(y^6\) is \(y^2\)

OpenStudy (anonymous):

it is supposed to be the cubed root of all of it together it's one problem

OpenStudy (anonymous):

the only one that stays inside the radical a \(y\) reason as follows : 3 goes in to 4 one time with a remainder of 1, one \(y\) comes out, one \(y\) stays in

OpenStudy (anonymous):

yes i know

OpenStudy (anonymous):

\[\sqrt[2]{ab}=\sqrt[3]{a}\sqrt[3]{b}\] so you can take it apart

OpenStudy (anonymous):

no that is wrong final answer should be \[2xy^2z\sqrt[3]{z}\]